During the year, Sarah’s Pet Shop’s merchandise inventory decreased by $43000. If the company’s cost of goods sold for the year was $900000, purchases would have been
Answer #1

Let beginning inventory=100,000

Hence ending inventory=(100,000-43000)=$57000

Cost of goods sold=Beginning inventory+Purchases-Ending inventory

900000=100,000+Purchases-57000

Purchases=900000+57000-100,000

=$857000